A Complete Bibliography of Publications in Higher-Order Symbolic Computation

Nelson H. F. Beebe Center for Scientific Computing

University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 2.15

Title word cross-reference

5 [3]. λ [142]. p [8]. ρ [154]. | [238].

-Calculus [8, 142, 154].

1960s [42]. 1999 [226]. 1999/2000 [226].

2007/2008 [226]. 2009 [206, 250]. 2009/ 2010 [250].

abortive [171]. Abstract [52, 70, 118, 138, 151, 160, 214]. Abstraction [40]. access [174]. accounting [178]. Accumulating [111]. accumulative [242]. Act [43]. Actor [11, 90]. AD [188].

Adaptation [110]. Adapting [189]. adaptive [225]. addresses [163]. algebra [144]. Algebraic [118]. algebras [146]. Algorithm [25, 59, 124]. Algorithmic [3]. Alias [124]. ambient [190]. Analysis [79, 104, 118, 124, 130, 151, 156, 213, 220, 230]. annotations [180, 181]. application [154, 156]. Applications [23]. Appreciation [123]. Approach [7, 35, 69, 109, 144, 198, 225, 245, 252]. approximation [230]. APTS [96]. Arithmetics [97]. arrays [227]. aspects [233]. Assignment [1, 106]. associated [248]. Augmented [108]. Automata [126]. Automatic [105, 185]. Automating [77]. axiomatics [147]. Axioms [87, 169].

Barrelled [83]. Based [29, 30, 72, 115, 130, 177, 179, 222, 251]. Be [20]. Beyond [23]. bidirectional [177]. binders [144]. Binding [10, 132, 147]. Bisimilarity [119]. BIT [131]. Bob [122, 123]. box [217]. BTI [184]. BTI-universal [184]. bytecode [213, 223].

C [236, 240]. Calculating [73]. Calculations [137, 210, 211]. Calculi [49]. Calculus [8, 16, 21, 58, 59, 80, 101, 119, 142, 149, 154, 159, 166, 190, 217]. Call [19, 20, 87, 150, 159, 192]. call-by-name [150, 159]. Call-by-push-value [150]. Call-by-Value [87, 150, 192]. Call/CC [19, 20]. calling [220]. calling-context [220]. calm [218]. Cambridge [42, 48]. Can [20]. Cartesian [222]. Certifying [33]. change [218]. check [231]. checking [231]. Christopher [39, 41, 44, 47, 48, 51]. circuit [193]. Circular [111, 232]. Class [18, 79, 215]. classical [142]. Code [33]. coinduction [234]. Collection [76]. Coloured [114]. Combinator [135]. combinatory [149]. Combining [32]. communication [251]. Compact [131]. Comparing [83, 100]. Compilation [33, 56, 192]. Compiler [35, 48, 113, 143]. Compiler-Compiler [35]. Compiling [21, 236, 240]. Completeness [80]. completing [245]. Composable [11]. compressed [243]. Computation [34, 35, 242]. Computational [97]. Computer [66, 205]. Computing [24, 211]. Concepts [37, 38, 46, 51]. concretization [140]. Connection [126]. consistency [231]. Constraint [110, 154]. constraints [155]. Construct [132]. Construction [62, 143]. Constructs [83]. Context [220]. Context-sensitive [220]. Contexts [24]. Continuation [19, 29, 30, 85, 236, 240]. Continuation-Based [29, 30]. Continuation-Passing [85, 236]. Continuations [18, 21, 53, 54, 86, 117, 171, 202, 203, 236, 240]. contracts [244]. Contributions [49]. Control [83, 169, 170, 174, 178]. Conversion [1, 84, 106]. correct [180, 181]. Correctness [89, 113, 161]. Corrigendum [106]. Cost [108]. CPL [42, 48]. CPS [1, 7, 23, 83, 84, 106, 169]. CPU [178]. cryptographic [156]. cryptoprotocol [153]. Cube [23]. Curry [135]. Cut [89]. cycles [253].

Data [32, 242, 243]. Declaring [109].

Decomposing [150]. Decomposition [77]. deep [246]. Definitional [13, 14]. Deforest [111]. defunctionalization [140]. delimited [170, 202, 203]. Dependent [81, 104, 153, 227, 231]. Derivation [124, 127, 232]. derivations [149]. Deriving [151]. description [193]. descriptions [231]. developing [177]. development [217]. Diagrams [90]. Difference [97]. Differencing [97]. Differentiation [66, 185]. Direct [8]. Directed [71]. Directly [196]. discrimination [222]. discrimination-based [222]. Distributed [76, 227]. distributive [147]. Divided [97]. Divided-Difference [97]. documents [177]. Domain [70, 138, 214]. Domains [49]. Double [83]. Double-Barrelled [83]. Dynamic [10, 56, 95, 139, 170, 179].

Early [21]. Eden [251]. editor [177]. Editorial [2, 12, 17, 26, 31, 36, 55, 60, 64, 68, 74, 78, 82, 88, 92, 98, 102, 107, 112, 116, 120, 129, 136, 141, 148, 152, 158, 168, 173, 182, 187, 191, 195, 200, 204, 219, 228, 241, 250]. effect [233]. effects [183]. Efficiency [61]. Efficient [127, 132, 174, 216]. Elimination [72]. embeddings [246]. Engineering [49]. Environment [59]. equality [215]. equational [176]. equations [249]. Erratum [240]. error [137, 229]. escape [151]. Euclidean [25]. Evaluation [21, 25, 34, 35, 63, 69, 71, 76, 108, 253]. events [236, 240]. Exceptions [20, 139]. Explaining [163]. Explicit [80, 145, 154, 163]. Expressible [193]. Expressing [149]. expressions [166]. Expressive [19]. Extended [16, 192]. Extension [135, 153].

F [215]. face [218]. Faithful [132]. families [248]. finite [137]. First [15, 18, 21, 142, 215]. First-Class [18, 215]. First-Order [21]. Fixed [135]. Fixed-Point [135]. Fixing [132]. Flattening [186]. flexible [174]. Flow [67, 86, 190]. Flow-sensitive [190]. Follow [43]. Foreword [37]. Formal [96, 99, 143]. Formalised [80]. Formalization [79]. forward [188]. forward-mode [188]. foundation [171, 202]. Foundations [99]. founded [180, 181]. framework [143, 188, 201]. FRP [218]. Full [54]. Functional [66, 91, 100, 108, 114, 155, 179, 188, 189, 192, 193, 225, 226, 233, 237, 238, 243, 245]. functions [242]. Fundamental [37, 38, 51]. Fusion [89, 232]. Futamura [71].

GADTs [215]. Garbage [76]. Generalization [5, 6, 89]. Generation [33, 105, 205, 223]. Generic [65, 155, 222, 225, 229]. Geometric [105]. Geometry [91]. Getting [208, 209]. Glossary [63]. Gödel [212]. gradual [216]. Graph [251–253]. Graph-based [251]. Graphical [165, 252]. Growing [27].

Handling [54]. Hard [43, 130]. hardware [248]. Haskell [198]. Heart [211]. hierarchy [169]. Higher [14, 57, 145, 189, 201, 232]. Higher-Order [14, 57, 145, 201, 232]. Hole [211]. Hop [224]. Hume [217]. Hybrid [69].

Imperative [175]. Imperative-program [175]. Implementation [18, 46, 76, 132, 172, 185]. Implementing [114, 134]. Implications [19]. implicitly [245]. Improved [124]. Improving [164, 229, 237]. incremental [252]. Incrementalization [61, 95]. Induction [49]. inequality [214]. Inference [59, 252]. Influence [41]. Information [67, 86]. initial [144]. inspired [205]. instrumented [175]. instrumented-interpreter [175]. Integrating [133]. interaction [244]. interconnects [231]. intermediate [186]. Interpretation [58, 118, 151]. Interpreter [16, 56, 175]. Interpreters [13, 14, 77]. Introduction [4, 5, 22, 61, 208]. inverse [242]. investigation [184]. iterator [212].

J [199]. Java [79, 178, 213, 229]. joins [222]. Jones [174, 184]. Jones-optimal [174]. Jumps [5, 6, 54].

Keeping [218]. Krivine [160–165].

Labels [5, 6, 208, 209]. Lambda

[16, 58, 80, 101, 159, 166]. Lambda-Calculus [80, 159, 166]. Lambda-I-Calculus [101]. Landin [5, 199, 205–208, 210]. Language [3, 27, 46, 49, 114, 153]. Languages [14, 37–39, 65, 70, 100, 192, 233, 245]. laws [147]. laziness [244]. lazy [163, 164, 179, 233, 237]. lean [215]. Least [125]. Letrec [132]. Level [133]. light [178]. light-weight [178]. Linear [58, 85, 86]. Linear-Time [58]. Linearity [212]. linguistic [246]. Lisp [194]. Lisp/CLOS [194]. Loading [79]. Local [84]. localised [213]. localization [233]. Logic [104, 108, 113, 126, 127, 153, 155, 174, 180, 181, 189]. Logic-Automata [126]. Logical [143, 230]. look [197]. Loops [84].

machine [159–161, 163–165, 178, 247].

machines [162, 166, 167].

Macrogeneration [40]. Magic [213]. Magic-sets [213]. Make [130]. Making [221]. Management [115]. Mathematical [54]. Maude [157]. Mechanising [142]. membership [176]. Memory [115]. messages [229]. meta [196]. meta-programming [196]. Metcast [134]. Microcontrollers [131]. minimalistic [197]. ML [9, 75]. Mnemonics [223]. mode [188]. Model [62, 67]. Models [11, 52]. modules [235]. Moessner [234]. MSR [153]. multi [224]. multi-tier [224]. multiple [242]. Multiprocessing [29, 30]. multiset [222]. My [45].

name [150, 159]. named [146]. names [245]. narrowing [156]. Nested [84, 198, 235]. Nesting [188]. Nets [114]. Network [75]. next [162]. Non [103]. Non-Standard [103]. Nondeterminism [127]. Normalization [101]. NSF [128].

Object [65, 183]. Object-Oriented [65]. octagon [138]. Offline [69]. Online [69]. OpenMusic [194]. Operational [99, 176]. operators [169, 197]. Optimal [28, 174]. optimality [184]. optimisation [218]. Optimization [113]. optimizations [201]. Optimizing [84]. Order [14, 21, 57, 142, 145, 189, 201, 232]. orientation [183]. Oriented [65]. Output [110]. Output-Constraint [110].

Paige [121–123]. Parallel [100]. Parameters [111]. Partial [21, 25, 34, 35, 63, 69, 71, 108]. Passing [85, 236, 240]. Path [94, 104, 235]. 2000 [226]. 2008 [226]. 2010 [250]. CC [19, 20]. CLOS [194]. PEPM2012 [241]. Performance [76, 100]. permutation [146]. permutations [142]. Personal [44, 206]. Peter [199, 205–207, 210]. Petri [114]. PLT [172]. Point [135]. Points [125]. Polymorphic [1, 28, 59, 106, 140]. Polynomial [242]. Polynomial-time [242]. Power [19, 118]. Powerful [20]. Pragmatic [109]. pragmatics [157]. pre [146]. precision [137]. Predicate [21]. Predicates [105]. presence [253]. principled [198]. printf [203]. Prize [207]. problem [203]. Process [34, 35, 239]. Processes [133]. products [222]. Program [32, 56, 70, 81, 96, 103, 109, 110, 174, 175, 178, 180, 181, 230]. programmable [177]. Programming [14, 37–39, 49, 95, 100, 114, 155, 194, 196, 198, 222, 225, 226, 239]. Programming-Language [49]. Programs [66, 67, 104, 108, 111, 127, 176, 179, 189, 232, 237, 243]. Projection [71]. Proof [77, 80, 89, 171, 234]. proof-theoretic [171]. proofs [215]. propagation [137]. Proposal [128]. Protocol [75]. protocols [156]. Proving [176]. Pruning [237]. Pseudo [147]. Pseudo-distributive [147]. Pure [58]. purely [233]. push [150].

Queries [94].

Racket [247]. randomized [247]. reachability [156]. Real [157]. Real-Time [157]. reasoning [218]. Recollection [44]. Recollections [41, 206]. Recursion [87, 192]. Recursive [132, 235]. redex [179]. reducing [160]. Reduction [127, 149, 253]. Reflections [50]. Reflective [65, 196]. Reflexive [125]. Region [115, 119]. Region-Based [115]. Regular [70, 94]. Reification [65]. Related [63]. Relations [125]. Relativizations [126]. relevant [221]. remembering [210]. Remembrances [122]. Reminiscences [47]. Report [3, 15]. representation [186]. Representations [28]. Representing [153]. Requirements [96]. Rerooting [76]. Research [93]. resolution [235]. restoration [139]. Retraction [7]. Retrospective [93, 115]. reuse [246]. Revised [3]. Revisited [13, 15, 25, 29, 34, 53, 167]. rewriting [144, 149, 153, 253]. Rid [208, 209]. Robert [121]. robust [217]. roundoff [137]. Rules [21, 232]. Run [33, 223]. Run-Time [33].

Safe [9, 223]. Safe-for-Space [9]. sanity [231]. Scala [246]. Scala-Virtualized [246]. Scenarios [109]. Scheme [3, 15, 16, 131, 132, 134, 155, 172]. scientist [205]. Scsh [133]. Second [71]. Secure [67, 86]. Selectors [130]. Self [58]. Self-Interpretation [58]. Semantic [11]. Semantics [54, 99, 103, 137, 157, 224]. sensitive [190, 220]. separation [239]. sequences [237]. Sequential [67]. server [172]. Servers [117]. Set [130]. sets [146, 213]. sharing [193]. sheaves [146]. Short [89]. Short-Cut [89]. Shortcut [232]. Side [233]. Side-effect [233]. Simply [80]. simulation [170, 248]. Sized [73]. Slicing [62, 103, 179]. Software [62, 217]. solutions [203]. Some [50]. Soundness [80]. Space [9, 216]. Space-efficient [216]. Specialisation [70, 174]. Specialization [32, 56, 109, 110, 127, 175]. specialized [239]. specializers [184]. Specification [90, 153, 215]. Specifications [96]. SSA [186]. Stack [75]. Staged [105, 239]. Standard [9, 75, 103]. State [139, 166, 167]. State-transition [166, 167]. Static [56, 95, 170, 231]. Storage [52]. Stories [210]. Strachey [39, 41, 44, 45, 47–51]. Strategies [18]. stricterness [221]. Strictly [20]. Strong [101]. Strongly [160]. Structural [249]. structured [177]. Study [211]. Style [8]. substitution [163]. Substitutions [80, 145]. Subtyping [28]. Symbolic [156, 222]. Syntactic [10, 40, 77]. syntax [145]. Synthesis [96, 248]. System [131, 212, 215, 217]. Systems [90, 144, 149, 190, 249].

Technique [65]. Temporal [113]. term [144]. Termination [81, 176]. terms [142]. testing [247]. TFP [250]. theorem [234]. theoretic [171, 202]. Theories [11, 77]. Theory [10, 142]. Threads [9, 133, 236, 240]. three [203]. tier [224]. Time [33, 58, 157, 223, 242]. Too [130]. tools [185].

Topics [63]. Totally [180, 181]. trace [183]. trails [179]. Transform [7]. transformation [175, 217, 252]. Transformational [124]. transformations [177, 178, 180, 181]. transition [166, 167]. Translation [101]. Translations [23]. traversals [225, 242]. Tree [70, 76]. Trends [226]. tuples [186]. Twice [19]. two [214]. Type [1, 59, 71, 72, 106, 190, 202, 223, 229, 239, 248, 252]. Type-Based [72]. Type-Directed [71]. Type-Inference [59]. type-safe [223]. Type-specialized [239]. type-theoretic [202]. Typed [80, 140, 245]. Types [28, 73, 81, 153, 183, 198, 212, 227, 231, 248, 249]. typing [203, 216].

Unchecked [20]. uncurrying [57, 201]. Understanding [39]. Unique [77]. Universal [94, 184]. unparsing [238]. Use [21, 172]. Useless [72]. Useless-Variable [72]. User [133]. User-Level [133]. Using [19, 84, 96, 111, 114, 139, 142, 156, 163, 225, 231].

Value [87, 150, 192]. Variable

[72, 144, 147, 214, 245]. Variadic [135]. variants [160]. Verification [81, 156]. verified [201]. Verilog [231]. versioning [139]. versus [226]. Very [131]. via [86, 95, 110, 174, 180, 181, 233]. Virtual [178, 247]. Virtualized [246]. Visual [194].

Weak [101]. Web [117, 172]. weight [178]. well [180, 181]. well-founded [180, 181]. who [205]. widening [197]. wire [231]. without [220]. Work [50]. Written [21].

Years [45].

References

Harper:1993:PTA

[1] Robert W. Harper and Mark Lillibridge. Polymorphic type assignment and CPS conversion. Lisp and Symbolic Computation, 6(3/4):361–380, November 1993. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic). See corrigendum [106].

Kessler:1998:E

[2] Bob Kessler and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 11(1):5, August 1998. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=11&issue=1&spage=5;   http:
//www.wkap.nl/oasis.htm/182522.

Abelson:1998:RRA

[3] H. Abelson, R. K. Dybvig, C. T. Haynes, G. J. Rozas, N. I. Adams IV, D. P. Friedman, E. Kohlbecker, G. L. Steele Jr., D. H. Bartley, R. Halstead, D. Oxley, G. J. Sussman, G. Brooks, C. Hanson, K. M. Pitman, and M. Wand. Revised ^{5} report on the algorithmic language Scheme. Higher-Order and Symbolic Computation, 11(1):7–105, August 1998. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=11&issue=1&spage=7;   http:
//www.wkap.nl/oasis.htm/168705.

Danvy:1998:I

[4] Olivier Danvy and Carolyn L. Talcott. Introduction. Higher-Order and Symbolic Computation, 11(2): 115–116, December 1998. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=11&issue=2&spage=115;   http:
//www.wkap.nl/oasis.htm/187565.

Thielecke:1998:ILG

[5] Hayo Thielecke. An introduction to Landin’s “A Generalization of Jumps and Labels”. Higher-Order and Symbolic Computation, 11(2): 117–123, December 1998. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=11&issue=2&spage=117;   http:
//www.wkap.nl/oasis.htm/187566.

Landin:1998:GJL

[6] Peter J. Landin. A generalization of jumps and labels. Higher-Order and Symbolic Computation, 11(2): 125–143, December 1998. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=11&issue=2&spage=125;   http:
//www.wkap.nl/oasis.htm/188501.

Kucan:1998:RAC

[7] Jakov Kucan. Retraction approach to CPS transform. Higher-Order and Symbolic Computation, 11(2): 145–175, December 1998. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&

volume=11&issue=2&spage=145;   http:
//www.wkap.nl/oasis.htm/187567.

Boudol:1998:CDS

[8] Gérard Boudol. The p-calculus in direct style. Higher-Order and Symbolic Computation, 11(2):177– 208, December 1998. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=11&issue=2&spage=177;   http:
//www.wkap.nl/oasis.htm/187568.

Biagioni:1998:SST

[9] Edoardo Biagioni, Ken Cline, Peter Lee, Chris Okasaki, and Chris Stone. Safe-for-space threads in Standard ML. Higher-Order and Symbolic Computation, 11(2):209– 225, December 1998. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=11&issue=2&spage=209;   http:
//www.wkap.nl/oasis.htm/187569.

Moreau:1998:STD

[10] Luc Moreau. A syntactic theory of dynamic binding. Higher-Order and Symbolic Computation, 11(3): 233–279, September 1998. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=11&issue=3&spage=233;   http:
//www.wkap.nl/oasis.htm/188358.

Talcott:1998:CSM

[11] Carolyn L. Talcott. Composable semantic models for actor theories. Higher-Order and Symbolic Computation, 11 (3):281–343, September 1998. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=11&issue=3&spage=281;   http:
                                        //www.wkap.nl/oasis.htm/188360.

Danvy:1998:E

[12] Olivier Danvy and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 11(4): 353–354, December 1998. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=11&issue=4&spage=353;   http:
                                        //www.wkap.nl/oasis.htm/193814.

Reynolds:1998:DIR

[13] John C. Reynolds. Definitional interpreters revisited. Higher-Order and Symbolic Computation, 11(4): 355–361, December 1998. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=11&issue=4&spage=355;   http:
                                        //www.wkap.nl/oasis.htm/193815.

Reynolds:1998:DIH

[14] John C. Reynolds. Definitional interpreters for higher-order programming languages. Higher-Order and Symbolic Computation, 11(4):363– 397, December 1998. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=11&issue=4&spage=363;   http:
//www.wkap.nl/oasis.htm/193816.

Sussman:1998:FRS

[15] Gerald Jay Sussman and Guy L. Steele Jr. The first report on Scheme revisited. Higher-Order and Symbolic Computation, 11(4): 399–404, December 1998. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=11&issue=4&spage=399;   http:
//www.wkap.nl/oasis.htm/193819.

Sussman:1998:SIE

[16] Gerald Jay Sussman and Guy L. Steele Jr. Scheme: An interpreter for extended lambda calculus. Higher-Order and Symbolic Computation, 11 (4):405–439, December 1998. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=11&issue=4&spage=405;   http:
//www.wkap.nl/oasis.htm/193825.

Danvy:1999:Ea

[17] Olivier Danvy and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 12(1):5, April 1999. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=12&issue=1&spage=5;   http:
//www.wkap.nl/oasis.htm/202966.

Clinger:1999:ISF

[18] William D. Clinger, Anne H. Hartheimer, and Eric M. Ost. Implementation strategies for first-class continuations. Higher-Order and Symbolic Computation, 12(1):7–45, April 1999. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=12&issue=1&spage=7;   http:
                                        //www.wkap.nl/oasis.htm/202967.

Thielecke:1999:UCT

[19] Hayo Thielecke. Using a continuation twice and its implications for the expressive power of call/cc. Higher-Order and Symbolic Computation, 12(1):47–73, April 1999. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=12&issue=1&spage=47;   http:
                                        //www.wkap.nl/oasis.htm/202968.

Lillibridge:1999:UEC

[20] Mark Lillibridge. Unchecked exceptions can be strictly more powerful than call/cc. Higher-Order and Symbolic Computation, 12(1):75–104, April 1999. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1388-3690&volume=12&issue=1&spage=
                                        75;   http://www.wkap.nl/oasis.
                                        htm/202969.

Sandewall:1999:EUC

[21] Erik Sandewall. An early use of continuations and partial evaluation for compiling rules written in firstorder predicate calculus. Higher-Order and Symbolic Computation, 12 (1):105–113, April 1999. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=12&issue=1&spage=105;   http:
//www.wkap.nl/oasis.htm/202970.

Danvy:1999:I

[22] Olivier Danvy and Carolyn Talcott. Introduction. Higher-Order and Symbolic Computation, 12(2): 123–124, September 1999. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=12&issue=2&spage=123;   http:
//www.wkap.nl/oasis.htm/234968.

Barthe:1999:CTA

[23] Gilles Barthe, John Hatcliff, and Morten Heine B. Sørensen. CPS translations and applications: The cube and beyond. Higher-Order and Symbolic Computation, 12(2): 125–170, September 1999. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=12&issue=2&spage=125;   http:
//www.wkap.nl/oasis.htm/234969.

Mason:1999:CC

[24] Ian A. Mason. Computing with contexts. Higher-Order and Symbolic Computation, 12(2):171–201, September 1999. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1388-3690&volume=12&issue=2&spage=
                                        171;   http://www.wkap.nl/oasis.
                                        htm/234970.

Lee:1999:PEE

[25] C. S. Lee. Partial evaluation of the Euclidean algorithm, revisited. Higher-Order and Symbolic Computation, 12 (2):203–212, September 1999. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=12&issue=2&spage=203;   http:
                                        //www.wkap.nl/oasis.htm/234971.

Danvy:1999:Eb

[26] Olivier Danvy and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 12(3):219, October 1999. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=12&issue=3&spage=219;   http:
                                        //www.wkap.nl/oasis.htm/237511.

Steele:1999:GL

[27] Guy L. Steele Jr. Growing a language. Higher-Order and Symbolic Computation, 12(3):221–236, October 1999. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=12&issue=3&spage=221;   http:
                                        //www.wkap.nl/oasis.htm/237513.

Aiken:1999:ORP

[28] Alexander Aiken, Edward L. Wimmers, and Jens Palsberg. Optimal representations of polymorphic types with subtyping. Higher-Order and Symbolic Computation, 12(3): 237–282, October 1999. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=12&issue=3&spage=237;   http:
//www.wkap.nl/oasis.htm/237515.

Wand:1999:CBMa

[29] Mitchell Wand. Continuation-based multiprocessing revisited. Higher-Order and Symbolic Computation, 12 (3):283, October 1999. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=12&issue=3&spage=283;   http:
//www.wkap.nl/oasis.htm/237519.

Wand:1999:CBMb

[30] Mitchell Wand. Continuation-based multiprocessing. Higher-Order and Symbolic Computation, 12(3):285–299, October 1999. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=1388-3690&volume=12&
issue=3&spage=285;   http://www.
wkap.nl/oasis.htm/237518.

Danvy:1999:Ec

[31] Olivier Danvy and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 12(4): 307–308, December 1999. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1388-3690&
                                        volume=12&issue=4&spage=307;   http:
                                        //www.wkap.nl/oasis.htm/247760.

Chirokoff:1999:CPD

[32] Sandrine Chirokoff, Charles Consel, and Renaud Marlet. Combining program and data specialization. Higher-Order and Symbolic Computation, 12 (4):309–335, December 1999. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=12&issue=4&spage=309;   http:
                                        //www.wkap.nl/oasis.htm/247762.

Hornof:1999:CCR

[33] Luke Hornof and Trevor Jim. Certifying compilation and run-time code generation. Higher-Order and Symbolic Computation, 12(4):337– 375, December 1999. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=12&issue=4&spage=337;   http:
                                        //www.wkap.nl/oasis.htm/247764.

Futamura:1999:PECa

[34] Yoshihiko Futamura. Partial evaluation of computation process, revisited. Higher-Order and Symbolic Computation, 12(4):377–380, December 1999. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1388-3690&volume=12&issue=4&spage=
                                        377;   http://www.wkap.nl/oasis.
                                        htm/247765.

Futamura:1999:PECb

[35] Yoshihiko Futamura. Partial evaluation of computation process — an approach to a compiler-compiler. Higher-Order and Symbolic Computation, 12 (4):381–391, December 1999. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=12&issue=4&spage=381;   http:
//www.wkap.nl/oasis.htm/247768.

Danvy:2000:Ea

[36] Olivier Danvy and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 13(1–2):5–6, April 2000. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=13&issue=1&spage=5;   http:
//www.wkap.nl/oasis.htm/257991.

Mosses:2000:FFC

[37] Peter D. Mosses. A foreword to “Fundamental Concepts in Programming Languages”. Higher-Order and Symbolic Computation, 13 (1–2):7–9, April 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=13&issue=1&spage=7;   http:
//www.wkap.nl/oasis.htm/257992.

Strachey:2000:FCP

[38] Christopher Strachey. Fundamental concepts in programming languages. Higher-Order and Symbolic Computation, 13(1–2):11–49, April 2000. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=13&issue=1&spage=11;   http:
                                        //www.wkap.nl/oasis.htm/257993.

Burstall:2000:CSU

[39] Rod Burstall. Christopher Strachey — understanding programming languages. Higher-Order and Symbolic Computation, 13(1–2):51–55, April 2000. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=13&issue=1&spage=51;   http:
                                        //www.wkap.nl/oasis.htm/258012.

Dybvig:2000:MSA

[40] R. Kent Dybvig. From macrogeneration to syntactic abstraction. Higher-Order and Symbolic Computation, 13 (1–2):57–63, April 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=13&issue=1&spage=57;   http:
                                        //www.wkap.nl/oasis.htm/258015.

Gordon:2000:CSR

[41] Mike Gordon. Christopher Strachey: Recollections of his influence. Higher-Order and Symbolic Computation, 13 (1–2):65–67, April 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=13&issue=1&spage=65;   http:
                                        //www.wkap.nl/oasis.htm/258017.

Hartley:2000:CC

[42] David Hartley. Cambridge and CPL in the 1960s. Higher-Order and Symbolic Computation, 13(1– 2):69–70, April 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=13&issue=1&spage=69;   http:
//www.wkap.nl/oasis.htm/258018.

Hoare:2000:HAF

[43] C. A. R. Hoare. A hard act to follow. Higher-Order and Symbolic Computation, 13(1–2):71–72, April 2000. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=13&issue=1&spage=71;   http:
//www.wkap.nl/oasis.htm/258019.

Jackson:2000:CSP

[44] Michael Jackson. Christopher Strachey: a personal recollection. Higher-Order and Symbolic Computation, 13 (1–2):73–74, April 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=13&issue=1&spage=73;   http:
//www.wkap.nl/oasis.htm/258020.

Landin:2000:MYS

[45] Peter J. Landin. My years with Strachey. Higher-Order and Symbolic Computation, 13(1–2):75–76, April 2000. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1388-3690&
                                        volume=13&issue=1&spage=75;   http:
                                        //www.wkap.nl/oasis.htm/258021.

Milne:2000:LCI

[46] Robert Milne. From language concepts to implementation concepts. Higher-Order and Symbolic Computation, 13 (1–2):77–81, April 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=13&issue=1&spage=77;   http:
                                        //www.wkap.nl/oasis.htm/258022.

Penrose:2000:RCS

[47] Roger Penrose. Reminiscences of Christopher Strachey. Higher-Order and Symbolic Computation, 13(1– 2):83–84, April 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=13&issue=1&spage=83;   http:
                                        //www.wkap.nl/oasis.htm/258023.

Richards:2000:CSC

[48] Martin Richards. Christopher Strachey and the Cambridge CPL compiler. Higher-Order and Symbolic Computation, 13(1–2):85–88, April 2000. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=13&issue=1&spage=85;   http:
                                        //www.wkap.nl/oasis.htm/258024.

Schmidt:2000:IDC

[49] David A. Schmidt. Induction, domains, calculi: Strachey’s contributions to programming-language engineering. Higher-Order and Symbolic Computation, 13(1–2):89–101, April 2000. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=13&issue=1&spage=89;   http:
//www.wkap.nl/oasis.htm/258025.

Scott:2000:SRS

[50] Dana Scott. Some reflections on Strachey and his work. Higher-Order and Symbolic Computation, 13 (1–2):103–114, April 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=13&issue=1&spage=103;   http:
//www.wkap.nl/oasis.htm/258026.

Stoy:2000:CSF

[51] Joe Stoy. Christopher Strachey and fundamental concepts. Higher-Order and Symbolic Computation, 13 (1–2):115–117, April 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=13&issue=1&spage=115;   http:
//www.wkap.nl/oasis.htm/258027.

Tennent:2000:AMS

[52] Robert D. Tennent and Dan R. Ghica. Abstract models of storage. Higher-Order and Symbolic Computation, 13 (1–2):119–129, April 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&

                                        volume=13&issue=1&spage=119;   http:
                                        //www.wkap.nl/oasis.htm/258028.

Wadsworth:2000:CR

[53] Christopher P. Wadsworth. Continuations revisited. Higher-Order and Symbolic Computation, 13(1– 2):131–133, April 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=13&issue=1&spage=131;   http:
                                        //www.wkap.nl/oasis.htm/258029.

Strachey:2000:CMS

[54] Christopher Strachey and Christopher P. Wadsworth. Continuations: a mathematical semantics for handling full jumps. Higher-Order and Symbolic Computation, 13(1– 2):135–152, April 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=13&issue=1&spage=135;   http:
                                        //www.wkap.nl/oasis.htm/258030.

Danvy:2000:Eb

[55] Olivier Danvy and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 13(3): 159–160, September 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=13&issue=3&spage=159;   http:
                                        //www.wkap.nl/oasis.htm/272138.

Thibault:2000:SDP

[56] Scott Thibault, Charles Consel, Julia L. Lawall, Renaud Marlet, and Gilles Muller. Static and dynamic program compilation by interpreter specialization. Higher-Order and Symbolic Computation, 13(3): 161–178, September 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=13&issue=3&spage=161;   http:
//www.wkap.nl/oasis.htm/272057.

Hannan:2000:HOU

[57] John Hannan and Patrick Hicks. Higher-order UnCurrying. Higher-Order and Symbolic Computation, 13 (3):179–216, September 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=13&issue=3&spage=179;   http:
//www.wkap.nl/oasis.htm/272058.

Mogensen:2000:LTS

[58] Torben Æ. Mogensen. Lineartime self-interpretation of the pure lambda calculus. Higher-Order and Symbolic Computation, 13(3):217– 237, September 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=13&issue=3&spage=217;   http:
//www.wkap.nl/oasis.htm/272059.

Nishizaki:2000:PEC

[59] Shin-Ya Nishizaki. A polymorphic environment calculus and its typeinference algorithm. Higher-Order and Symbolic Computation, 13(3): 239–278, September 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=13&issue=3&spage=239;   http:
//www.wkap.nl/oasis.htm/272061.

Danvy:2000:Ec

[60] Olivier Danvy and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 13(4): 287–288, December 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=13&issue=4&spage=287;   http:
//www.wkap.nl/oasis.htm/278950.

Liu:2000:EII

[61] Yanhong A. Liu. Efficiency by incrementalization: An introduction. Higher-Order and Symbolic Computation, 13 (4):289–313, December 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=13&issue=4&spage=289;   http:
//www.wkap.nl/oasis.htm/278951.

Hatcliff:2000:SSM

[62] John Hatcliff, Matthew B. Dwyer, and Hongjun Zheng. Slicing software for model construction. Higher-Order and Symbolic Computation, 13 (4):315–353, December 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=1388-3690&
volume=13&issue=4&spage=315;   http:
//www.wkap.nl/oasis.htm/278952.

Mogensen:2000:GPE

[63] Torben Æ. Mogensen. Glossary for partial evaluation and related topics. Higher-Order and Symbolic Computation, 13(4):355–368, December 2000. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=13&issue=4&spage=
355;   http://www.wkap.nl/oasis.
htm/278953.

Danvy:2001:Ea

[64] Olivier Danvy and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 14(1):5, March 2001. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=14&issue=1&spage=5;   http:
//www.wkap.nl/oasis.htm/323056.

Douence:2001:GRT

[65] Rémi Douence and Mario Südholt. A generic reification technique for objectoriented reflective languages. Higher-Order and Symbolic Computation, 14 (1):7–34, March 2001. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=14&issue=1&spage=7;   http:
//www.wkap.nl/oasis.htm/323057.

Karczmarczuk:2001:FDC

[66] Jerzy Karczmarczuk. Functional differentiation of computer programs. Higher-Order and Symbolic Computation, 14(1):35–57, March 2001. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=14&issue=1&spage=35;   http:
//www.wkap.nl/oasis.htm/323058.

Sabelfeld:2001:PMS

[67] Andrei Sabelfeld and David Sands. A per model of secure information flow in sequential programs. Higher-Order and Symbolic Computation, 14 (1):59–91, March 2001. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=14&issue=1&spage=59;   http:
//www.wkap.nl/oasis.htm/323060.

Danvy:2001:Eb

[68] Olivier Danvy and Julia L. Lawall. Editorial. Higher-Order and Symbolic Computation, 14(2–3):99–100, September 2001. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=14&issue=2&spage=
99;   http://www.wkap.nl/oasis.
htm/323060.

Sumii:2001:HAO

[69] Eijiro Sumii and Naoki Kobayashi. A hybrid approach to online and offline partial evaluation. Higher-Order and Symbolic Computation, 14(2–3): 101–142, September 2001. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=14&issue=2&spage=101;   http:
//www.wkap.nl/oasis.htm/323060.

Gallagher:2001:RTL

[70] John P. Gallagher and Julio C. Peralta. Regular tree languages as an abstract domain in program specialisation. Higher-Order and Symbolic Computation, 14(2–3):143– 172, September 2001. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=14&issue=2&spage=143;   http:
//www.wkap.nl/oasis.htm/323060.

Grobauer:2001:SFP

[71] Bernd Grobauer and Zhe Yang. The second Futamura projection for typedirected partial evaluation. Higher-Order and Symbolic Computation, 14 (2–3):173–219, September 2001. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=14&issue=2&spage=173;   http:
//www.wkap.nl/oasis.htm/323060.

Kobayashi:2001:TBU

[72] Naoki Kobayashi. Type-based uselessvariable elimination. Higher-Order and Symbolic Computation, 14(2–3): 221–260, September 2001. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=14&issue=2&spage=221;   http:
//www.wkap.nl/oasis.htm/323060.

Chin:2001:CST

[73] Wei-Ngan Chin and Siau-Cheng Khoo. Calculating sized types. Higher-Order and Symbolic Computation, 14(2–3): 261–300, September 2001. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=14&issue=2&spage=261;   http:
                                        //www.wkap.nl/oasis.htm/323060.

Danvy:2001:Ec

[74] Olivier Danvy, Takayasu Ito, and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 14 (4):307, December 2001. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:/

                                        /ipsapp007.lwwonline.com/content/
                                        getfile/4979/18/1/abstract.htm;
                                        http://ipsapp007.lwwonline.com/
                                        content/getfile/4979/18/1/fulltext.
                                        pdf;   http://www.springerlink.
                                        com/openurl.asp?genre=article&issn=
                                        1388-3690&volume=14&issue=4&spage=
                                        307.

Biagioni:2001:NPS

[75] Edoardo Biagioni, Robert Harper, and Peter Lee. A network protocol stack in Standard ML. Higher-Order and Symbolic Computation, 14 (4):309–356, December 2001. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:/

                                        /ipsapp007.lwwonline.com/content/
                                        getfile/4979/18/2/abstract.htm;
                                        http://ipsapp007.lwwonline.com/
                                        content/getfile/4979/18/2/fulltext.
                                        pdf;   http://www.springerlink.
                                        com/openurl.asp?genre=article&issn=

1388-3690&volume=14&issue=4&spage=
309.

Moreau:2001:TRD

[76] Luc Moreau. Tree rerooting in distributed garbage collection: Implementation and performance evaluation. Higher-Order and Symbolic Computation, 14(4):357–386, December 2001. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL

http://ipsapp007.lwwonline.com/
content/getfile/4979/18/3/abstract.
htm;   http://ipsapp007.lwwonline.
com/content/getfile/4979/18/3/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
1388-3690&volume=14&issue=4&spage=
357.

Xiao:2001:STI

[77] Yong Xiao, Amr Sabry, and Zena M. Ariola. From syntactic theories to interpreters: Automating the proof of unique decomposition. Higher-Order and Symbolic Computation, 14 (4):387–409, December 2001. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:/

/ipsapp007.lwwonline.com/content/
getfile/4979/18/4/abstract.htm;
http://ipsapp007.lwwonline.com/
content/getfile/4979/18/4/fulltext.
pdf;   http://www.springerlink.
com/openurl.asp?genre=article&issn=
1388-3690&volume=14&issue=4&spage=
387.

Danvy:2002:Ea

[78] Olivier Danvy, Takayasu Ito, and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 15 (1):5, March 2002. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/content/getfile/
                                        4979/19/1/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/4979/19/1/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=15&issue=1&spage=5.

Tozawa:2002:FAC

[79] Akihiko Tozawa and Masami Hagiya. Formalization and analysis of class loading in Java. Higher-Order and Symbolic Computation, 15(1):7–55, March 2002. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/content/getfile/
                                        4979/19/2/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/4979/19/2/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=15&issue=1&spage=7.

Coquand:2002:FPS

[80] Catarina Coquand. A formalised proof of the soundness and completeness of a simply typed lambda-calculus with explicit substitutions. Higher-Order and Symbolic Computation, 15(1):57– 90, March 2002. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/content/getfile/
                                        4979/19/3/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/4979/19/3/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=15&issue=1&spage=57.

Xi:2002:DTP

[81] Hongwei Xi. Dependent types for program termination verification. Higher-Order and Symbolic Computation, 15(1): 91–131, March 2002. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/19/4/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/19/4/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=15&issue=1&spage=91.

Danvy:2002:Eb

[82] Olivier Danvy and Amr Sabry. Editorial. Higher-Order and Symbolic Computation, 15(2–3):139–140, September 2002. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/20/1/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/20/1/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=15&issue=2&spage=139.

Thielecke:2002:CCC

[83] Hayo Thielecke. Comparing control constructs by double-barrelled CPS. Higher-Order and Symbolic Computation, 15(2–3):141–160, September 2002. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/20/2/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/20/2/fulltext.pdf;

                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=15&issue=2&spage=141.

Reppy:2002:ONL

[84] John Reppy. Optimizing nested loops using local CPS conversion. Higher-Order and Symbolic Computation, 15(2–3):161–180, September 2002. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/content/getfile/
                                        4979/20/3/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/4979/20/3/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=15&issue=2&spage=161.

Berdine:2002:LCP

[85] Josh Berdine, Peter O’Hearn, Uday Reddy, and Hayo Thielecke. Linear continuation-passing. Higher-Order and Symbolic Computation, 15(2–3): 181–208, September 2002. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http://

                                        ipsapp008.kluweronline.com/content/
                                        getfile/4979/20/4/abstract.htm;
                                         http://ipsapp008.kluweronline.
                                        com/content/getfile/4979/20/4/fulltext.
                                        pdf;   http://www.springerlink.
                                        com/openurl.asp?genre=article&issn=
                                        1388-3690&volume=15&issue=2&spage=
                                        181.

Zdancewic:2002:SIF

[86] Steve Zdancewic and Andrew C. Myers. Secure information flow via linear continuations. Higher-Order and Symbolic Computation, 15(2–3):209–234, September 2002. CODEN LSCOEX.

ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/20/5/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/20/5/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=15&issue=2&spage=209.

Hasegawa:2002:ARC

[87] Masahito Hasegawa and Yoshihiko Kakutani. Axioms for recursion in call-by-value. Higher-Order and Symbolic Computation, 15(2–3):235–264, September 2002. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/20/6/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/20/6/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=15&issue=2&spage=235.

Danvy:2002:Ec

[88] Olivier Danvy and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 15(4):271, December 2002. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/21/1/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/21/1/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=15&issue=4&spage=271.

Johann:2002:GSC

[89] Patricia Johann. A generalization of short-cut fusion and its correctness proof. Higher-Order and Symbolic Computation, 15(4):273–300, December 2002. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/21/2/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/21/2/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=15&issue=4&spage=273.

Smith:2002:SDA

[90] Scott F. Smith and Carolyn L. Talcott. Specification diagrams for actor systems. Higher-Order and Symbolic Computation, 15(4):301–348, December 2002. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/21/3/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/21/3/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=15&issue=4&spage=301.

Henderson:2002:FG

[91] Peter Henderson. Functional geometry. Higher-Order and Symbolic Computation, 15(4):349–365, December 2002. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/21/4/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/21/4/fulltext.pdf;
http://www.springerlink.com/openurl.

asp?genre=article&issn=1388-3690&
volume=15&issue=4&spage=349.

Danvy:2003:Ea

[92] Olivier Danvy, Fritz Henglein, Harry Mairson, and Alberto Pettorossi. Editorial. Higher-Order and Symbolic Computation, 16(1–2):5–6, March–June 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/22/1/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/22/1/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=16&issue=1&spage=5.

Paige:2003:RR

[93] Bob Paige. Research retrospective. Higher-Order and Symbolic Computation, 16(1–2):7–13, March–June 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/22/2/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/22/2/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=16&issue=1&spage=7.

deMoor:2003:URP

[94] Oege de Moor, David Lacey, and Eric Van Wyk. Universal regular path queries. Higher-Order and Symbolic Computation, 16(1–2):15–35, March– June 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/

                                        4979/22/3/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/4979/22/3/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=16&issue=1&spage=15.

Liu:2003:DPS

[95] Yanhong A. Liu and Scott D. Stoller. Dynamic programming via static incrementalization. Higher-Order and Symbolic Computation, 16(1–2):37–62, March–June 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/content/getfile/
                                        4979/22/4/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/4979/22/4/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=16&issue=1&spage=37.

Leonard:2003:PSF

[96] Elizabeth I. Leonard and Constance L. Heitmeyer. Program synthesis from formal requirements specifications using APTS. Higher-Order and Symbolic Computation, 16(1–2):63–92, March– June 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/content/getfile/
                                        4979/22/5/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/4979/22/5/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=16&issue=1&spage=63.

Reps:2003:CDD

[97] Thomas W. Reps and Louis B. Rall. Computational divided differencing and divided-difference arithmetics. Higher-Order and Symbolic Computation, 16(1–2):93–149, March– June 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/22/6/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/22/6/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=16&issue=1&spage=93.

Basin:2003:E

[98] David Basin, Olivier Danvy, Julian Padget, and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 16(3):159, September 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/23/1/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/23/1/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=16&issue=3&spage=159.

Ford:2003:FFO

[99] Jonathan Ford and Ian A. Mason. Formal foundations of operational semantics. Higher-Order and Symbolic Computation, 16(3):161–202, September 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/23/2/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/23/2/fulltext.pdf;
http://www.springerlink.com/openurl.

                                        asp?genre=article&issn=1388-3690&
                                        volume=16&issue=3&spage=161.

Loidl:2003:CPF

[100] H.-W. Loidl, F. Rubio, N. Scaife, K. Hammond, S. Horiguchi, U. Klusik, R. Loogen, G. J. Michaelson, R. Peña, S. Priebe, Á. J. Rebón, and P. W. Trinder. Comparing parallel functional languages: Programming and performance. Higher-Order and Symbolic Computation, 16(3):203–251, September 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/content/getfile/
                                        4979/23/3/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/4979/23/3/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=16&issue=3&spage=203.

Gortz:2003:SNW

[101] Inge Li Gørtz, Signe Reuss, and Morten Heine Sørensen. Strong normalization from weak normalization by translation into the lambda-I-calculus. Higher-Order and Symbolic Computation, 16(3):253–285, September 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/content/getfile/
                                        4979/23/4/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/4979/23/4/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=16&issue=3&spage=253.

Danvy:2003:Eb

[102] Olivier Danvy and Peter Thiemann.

Editorial. Higher-Order and Symbolic Computation, 16(4):295, December 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/24/1/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/24/1/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=16&issue=4&spage=295.

Giacobazzi:2003:NSS

[103] Roberto Giacobazzi and Isabella Mastroeni. Non-standard semantics for program slicing. Higher-Order and Symbolic Computation, 16(4):297–339, December 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/24/2/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/24/2/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=16&issue=4&spage=297.

Lu:2003:PDA

[104] Lunjin Lu. Path dependent analysis of logic programs. Higher-Order and Symbolic Computation, 16(4):341–377, December 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/content/getfile/
4979/24/3/abstract.htm;   http://
ipsapp008.kluweronline.com/content/
getfile/4979/24/3/fulltext.pdf;
http://www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=16&issue=4&spage=341.

Nanevski:2003:AGS

[105] Aleksandar Nanevski, Guy Blelloch, and Robert Harper. Automatic generation of staged geometric predicates. Higher-Order and Symbolic Computation, 16(4):379–400, December 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/content/getfile/
                                        4979/24/4/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/4979/24/4/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=16&issue=4&spage=379.

Harper:2003:CPT

[106] Robert Harper and Mark Lillibridge. Corrigendum: Polymorphic type assignment and CPS conversion. Higher-Order and Symbolic Computation, 16(4):401, December 2003. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/content/getfile/
                                        4979/24/5/abstract.htm;   http://
                                        ipsapp008.kluweronline.com/content/
                                        getfile/4979/24/5/fulltext.pdf;
                                        http://www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=16&issue=4&spage=401. See

[1].

Chin:2004:E

[107] Wei-Ngan Chin, Olivier Danvy, and Peter Thiemann. Editorial. Higher-Order and Symbolic Computation, 17 (1–2):5–6, March 2004. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=1388-3690&
volume=17&issue=1&spage=5.

Vidal:2004:CAP

[108] Germán Vidal. Cost–augmented partial evaluation of functional logic programs. Higher-Order and Symbolic Computation, 17(1–2):7–46, March 2004. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=17&issue=1&spage=7.

LeMeur:2004:SSP

[109] Anne-Françoise Le Meur, Julia L.

Lawall, and Charles Consel. Specialization scenarios: a pragmatic approach to declaring program specialization. Higher-Order and Symbolic Computation, 17(1–2):47–92, March 2004. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=17&issue=1&spage=47.

Khoo:2004:PAO

[110] Siau-Cheng Khoo and Kun Shi. Program adaptation via output-constraint specialization. Higher-Order and Symbolic Computation, 17(1–2):93–128, March 2004. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=
article&issn=1388-3690&volume=17&
issue=1&spage=93.

Voigtlander:2004:UCP

[111] Janis Voigtländer. Using circular programs to deforest in accumulating parameters. Higher-Order and Symbolic Computation, 17(1–2): 129–163, March 2004. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=17&issue=1&spage=129.

Basin:2004:E

[112] David Basin, Olivier Danvy, and Robert Harper. Editorial. Higher-Order and Symbolic Computation, 17 (3):171, September 2004. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=17&issue=3&spage=171.

Lacey:2004:COC

[113] David Lacey, Neil D. Jones, Eric Van Wyk, and Carl Christian Frederiksen. Compiler optimization correctness by temporal logic. Higher-Order and Symbolic Computation, 17 (3):173–206, September 2004. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=17&issue=3&spage=173.

Kristensen:2004:ICP

[114] Lars Michael Kristensen and Søren Christensen. Implementing coloured Petri nets using a functional programming language. Higher-Order and Symbolic Computation, 17(3): 207–243, September 2004. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=1388-3690&
volume=17&issue=3&spage=207.

Tofte:2004:RRB

[115] Mads Tofte, Lars Birkedal, Martin Elsman, and Niels Hallenberg. A retrospective on region-based memory management. Higher-Order and Symbolic Computation, 17(3): 245–265, September 2004. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=17&issue=3&spage=245.

Danvy:2004:E

[116] Olivier Danvy and Ian Mason. Editorial. Higher-Order and Symbolic Computation, 17(4):275, December 2004. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=17&issue=4&spage=275.

Queinnec:2004:CWS

[117] Christian Queinnec. Continuations and Web servers. Higher-Order and Symbolic Computation, 17(4): 277–295, December 2004. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=17&issue=4&spage=277.

Mastroeni:2004:APA

[118] Isabella Mastroeni. Algebraic power analysis by abstract interpretation. Higher-Order and Symbolic Computation, 17(4):297–345, December 2004.

CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=17&issue=4&spage=
297.

Helsen:2004:BRC

[119] Simon Helsen. Bisimilarity for the region calculus. Higher-Order and Symbolic Computation, 17(4): 347–394, December 2004. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=17&issue=4&spage=347.

Danvy:2005:E

[120] Olivier Danvy, Fritz Henglein, Harry Mairson, and Alberto Pettorossi. Editorial. Higher-Order and Symbolic Computation, 18(1–2):5–6, June 2005. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=18&issue=1&spage=5.

Mairson:2005:RP

[121] Harry Mairson. Robert Paige (1947– 1999). Higher-Order and Symbolic Computation, 18(1–2):7–8, June 2005. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=18&issue=1&spage=7.

Siegel:2005:RBP

[122] Alan Siegel. Remembrances of Bob Paige. Higher-Order and Symbolic Computation, 18(1–2):9–11, June 2005. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=18&issue=1&spage=9.

Davis:2005:ABP

[123] Martin Davis. An appreciation of Bob Paige. Higher-Order and Symbolic Computation, 18(1–2):13, June 2005. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=18&issue=1&spage=13.

Goyal:2005:TDI

[124] Deepak Goyal. Transformational derivation of an improved alias analysis algorithm. Higher-Order and Symbolic Computation, 18(1–2):15–49, June 2005. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=18&issue=1&spage=15.

Desharnais:2005:LRP

[125] Jules Desharnais and Bernhard Möller.

Least reflexive points of relations. Higher-Order and Symbolic Computation, 18(1–2):51–77, June 2005. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=18&issue=1&spage=51.

Klarlund:2005:RLA

[126] Nils Klarlund. Relativizations for the logic-automata connection. Higher-Order and Symbolic Computation, 18 (1–2):79–120, June 2005. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=18&issue=1&spage=79.

Pettorossi:2005:DEL

[127] Alberto Pettorossi, Maurizio Proietti, and Sophie Renault. Derivation of efficient logic programs by specialization and reduction of nondeterminism. Higher-Order and Symbolic Computation, 18(1–2):121–210, June 2005. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=18&issue=1&spage=121.

Paige:2005:NP

[128] Robert Paige. An NSF proposal. Higher-Order and Symbolic Computation, 18 (1–2):211–235, June 2005. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=18&issue=1&spage=211.

Felleisen:2005:E

[129] Matthias Felleisen, Julia Lawall, Manuel Serrano, and Olin Shivers. Editorial. Higher-Order and Symbolic Computation, 18(3–4):243–244, December 2005. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1388-3690&volume=18&issue=3&spage=
                                        243.

Meunier:2005:SMS

[130] Philippe Meunier, Robert Bruce Findler, Paul Steckler, and Mitchell Wand. Selectors make set–based analysis too hard. Higher-Order and Symbolic Computation, 18(3–4): 245–269, December 2005. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=18&issue=3&spage=245.

Dube:2005:BVC

[131] Danny Dubé and Marc Feeley. BIT:

a very compact scheme system for microcontrollers. Higher-Order and Symbolic Computation, 18(3–4):271– 298, December 2005. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=18&issue=3&spage=271.

Waddell:2005:FLF

[132] Oscar Waddell, Dipanwita Sarkar, and R. Kent Dybvig. Fixing Letrec: a faithful yet efficient implementation of Scheme’s recursive binding construct. Higher-Order and Symbolic Computation, 18(3–4): 299–326, December 2005. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=18&issue=3&spage=299.

Gasbichler:2005:IUL

[133] Martin Gasbichler and Michael Sperber. Integrating user-level threads with processes in Scsh. Higher-Order and Symbolic Computation, 18(3–4): 327–354, December 2005. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=18&issue=3&spage=327.

Kiselyov:2005:IMS

[134] Oleg Kiselyov. Implementing Metcast in Scheme. Higher-Order and Symbolic Computation, 18(3–4):355– 370, December 2005. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=18&issue=3&spage=355.

Goldberg:2005:VEC

[135] Mayer Goldberg. A variadic extension of Curry’s fixed-point combinator. Higher-Order and Symbolic Computation, 18 (3–4):371–388, December 2005. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=18&issue=3&spage=371.

Danvy:2006:Ea

[136] Olivier Danvy, Oege de Moor, Julian Padget, and Peter Thiemann. Editorial. Higher-Order and Symbolic Computation, 19(1):5, March 2006. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=19&issue=1&spage=5.

Martel:2006:SRE

[137] Matthieu Martel. Semantics of roundoff error propagation in finite precision calculations. Higher-Order and Symbolic Computation, 19 (1):7–30, March 2006. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=19&issue=1&spage=7.

Mine:2006:OAD

[138] Antoine Miné. The octagon abstract domain. Higher-Order and Symbolic Computation, 19(1):31–100, March 2006. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=19&issue=1&spage=
31.

Nandivada:2006:DSR

[139] V. Krishna Nandivada and Suresh Jagannathan. Dynamic state restoration using versioning exceptions. Higher-Order and Symbolic Computation, 19 (1):101–124, March 2006. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=19&issue=1&spage=101.

Pottier:2006:PTD

[140] François Pottier and Nadji Gauthier.

Polymorphic typed defunctionalization and concretization. Higher-Order and Symbolic Computation, 19 (1):125–162, March 2006. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=19&issue=1&spage=125.

Honsell:2006:E

[141] Furio Honsell and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 19(2–3):167– 168, September 2006. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=19&issue=2&spage=167.

Norrish:2006:MCU

[142] Michael Norrish. Mechanising λ-calculus using a classical first order theory of terms with permutations. Higher-Order and Symbolic Computation, 19(2–3): 169–195, September 2006. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=19&issue=2&spage=169.

Hickey:2006:FCC

[143] Jason Hickey and Aleksey Nogin. Formal compiler construction in a logical framework. Higher-Order and Symbolic Computation, 19(2–3):197– 230, September 2006. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=19&issue=2&spage=197.

Hamana:2006:IAA

[144] Makoto Hamana. An initial algebra approach to term rewriting systems with variable binders. Higher-Order and Symbolic Computation, 19 (2–3):231–262, September 2006. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=19&issue=2&spage=231.

Ghani:2006:ESH

[145] Neil Ghani, Tarmo Uustalu, and Makoto Hamana. Explicit substitutions and higher-order syntax. Higher-Order and Symbolic Computation, 19(2–3): 263–282, September 2006. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=19&issue=2&spage=263.

Gadducci:2006:APA

[146] Fabio Gadducci, Marino Miculan, and Ugo Montanari. About permutation algebras, (pre)sheaves and named sets. Higher-Order and Symbolic Computation, 19(2–3):283– 304, September 2006. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=19&issue=2&spage=283.

Tanaka:2006:PDL

[147] Miki Tanaka and John Power. Pseudodistributive laws and axiomatics for variable binding. Higher-Order and Symbolic Computation, 19(2–3):305– 337, September 2006. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=19&issue=2&spage=305.

Danvy:2006:Eb

[148] Olivier Danvy, Andrzej Filinski, Jean-Louis Giavitto, Andy King, Pierre-Etienne Moreau, and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 19(4): 343–344, December 2006. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=19&issue=4&spage=343.

Bertolissi:2006:ECR

[149] Clara Bertolissi, Horatiu Cirstea, and Claude Kirchner. Expressing combinatory reduction systems derivations in the rewriting calculus. Higher-Order and Symbolic Computation, 19 (4):345–376, December 2006. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=19&issue=4&spage=345.

Levy:2006:CPV

[150] Paul Blain Levy. Call-by-push-value: Decomposing call-by-value and call-by-name. Higher-Order and Symbolic Computation, 19(4): 377–414, December 2006. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=19&issue=4&spage=377.

Hill:2006:DEA

[151] Patricia M. Hill and Fausto Spoto. Deriving escape analysis by abstract interpretation. Higher-Order and Symbolic Computation, 19(4): 415–463, December 2006. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=19&issue=4&spage=415.

Marti-Oliet:2007:E

[152] Narciso Martí-Oliet, Grigore Rosu, and Carolyn Talcott. Editorial. Higher-Order and Symbolic Computation, 20 (1–2):1–2, June 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=20&issue=1&spage=1.

Cervesato:2007:RMC

[153] Iliano Cervesato and Mark-Oliver Stehr.

Representing the MSR cryptoprotocol specification language in an extension of rewriting logic with dependent types. Higher-Order and Symbolic Computation, 20(1–2):3–35, June 2007. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=20&issue=1&spage=3.

Cirstea:2007:CEC

[154] Horatiu Cirstea, Germain Faure, and Claude Kirchner. A ρ-calculus of explicit constraint application. Higher-Order and Symbolic Computation, 20 (1–2):37–72, June 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=20&issue=1&spage=37.

Fraguas:2007:NGS

[155] Francisco J. López Fraguas, Mario Rodríguez Artalejo, and Rafael del Vado Vírseda. A new generic scheme for functional logic programming with constraints. Higher-Order and Symbolic Computation, 20(1–2):73–122, June 2007. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=20&issue=1&spage=73.

Meseguer:2007:SRA

[156] José Meseguer and Prasanna Thati. Symbolic reachability analysis using narrowing and its application to verification of cryptographic protocols. Higher-Order and Symbolic Computation, 20 (1–2):123–160, June 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=20&issue=1&spage=123.

Olveczky:2007:SPR

[157] Peter Csaba Ölveczky and José Meseguer. Semantics and pragmatics of Real-Time Maude. Higher-Order and Symbolic Computation, 20 (1–2):161–196, June 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=20&issue=1&spage=161.

Danvy:2007:Ea

[158] Olivier Danvy. Editorial. Higher-Order and Symbolic Computation, 20 (3):197–198, September 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=20&issue=3&spage=197.

Krivine:2007:CNL

[159] Jean-Louis Krivine. A call-by-name lambda-calculus machine. Higher-Order and Symbolic Computation, 20 (3):199–207, September 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=20&issue=3&spage=199.

Cregut:2007:SRV

[160] Pierre Crégut. Strongly reducing variants of the Krivine abstract machine. Higher-Order and Symbolic Computation, 20(3):209– 230, September 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=20&issue=3&spage=209.

Wand:2007:CKM

[161] Mitchell Wand. On the correctness of the Krivine machine. Higher-Order and Symbolic Computation, 20 (3):231–235, September 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=20&issue=3&spage=231.

Douence:2007:NKM

[162] Rémi Douence and Pascal Fradet. The next 700 Krivine machines. Higher-Order and Symbolic Computation, 20 (3):237–255, September 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=20&issue=3&spage=237.

Lang:2007:ELK

[163] Frédéric Lang. Explaining the lazy Krivine machine using explicit substitution and addresses. Higher-Order and Symbolic Computation, 20 (3):257–270, September 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=20&issue=3&spage=257.

Friedman:2007:ILK

[164] Daniel P. Friedman, Abdulaziz Ghuloum, Jeremy G. Siek, and Onnie Lynn Winebarger. Improving the lazy Krivine machine. Higher-Order and Symbolic Computation, 20 (3):271–293, September 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=20&issue=3&spage=271.

Lippi:2007:GKM

[165] Sylvain Lippi. The graphical Krivine machine. Higher-Order and Symbolic Computation, 20(3):295– 318, September 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=20&issue=3&spage=295.

Schmidt:2007:STMa

[166] David A. Schmidt. State-transition machines for lambda-calculus expressions. Higher-Order and Symbolic Computation, 20(3):319–332, September 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=20&issue=3&spage=
319.

Schmidt:2007:STMb

[167] David A. Schmidt. State-transition machines, revisited. Higher-Order and Symbolic Computation, 20(3): 333–335, September 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=20&issue=3&spage=333.

Danvy:2007:Eb

[168] Olivier Danvy and Hayo Thielecke. Editorial. Higher-Order and Symbolic Computation, 20(4):337–338, December 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=20&issue=4&spage=
337.

Kameyama:2007:ACO

[169] Yukiyoshi Kameyama. Axioms for control operators in the CPS hierarchy. Higher-Order and Symbolic Computation, 20(4):339–369, December 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=

                                        1388-3690&volume=20&issue=4&spage=
                                        339.

Shan:2007:SSD

[170] Chung chieh Shan. A static simulation of dynamic delimited control. Higher-Order and Symbolic Computation, 20 (4):371–401, December 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=20&issue=4&spage=371.

Ariola:2007:PTF

[171] Zena M. Ariola, Hugo Herbelin, and Amr Sabry. A proof-theoretic foundation of abortive continuations. Higher-Order and Symbolic Computation, 20 (4):403–429, December 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=20&issue=4&spage=403.

Krishnamurthi:2007:IUP

[172] Shriram Krishnamurthi, Peter Walton Hopkins, Jay McCarthy, Paul T. Graunke, Greg Pettyjohn, and Matthias Felleisen. Implementation and use of the PLT scheme Web server. Higher-Order and Symbolic Computation, 20 (4):431–460, December 2007. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=20&issue=4&spage=431.

Lawall:2008:E

[173] Julia Lawall, Michael Leuschel, and Peter Sestoft. Editorial. Higher-Order and Symbolic Computation, 21 (1–2):1–3, June 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=21&issue=1&spage=1.

Barker:2008:EFA

[174] Steve Barker, Michael Leuschel, and Mauricio Varea. Efficient and flexible access control via Jones-optimal logic program specialisation. Higher-Order and Symbolic Computation, 21 (1–2):5–35, June 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=21&issue=1&spage=5.

Debois:2008:IPT

[175] Søren Debois. Imperative-program transformation by instrumented-interpreter specialization. Higher-Order and Symbolic Computation, 21(1–2):37–58, June 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=21&issue=1&spage=
37.

Duran:2008:POT

[176] Francisco Durán, Salvador Lucas, Claude Marché, José Meseguer, and Xavier Urbain. Proving operational termination of membership equational programs. Higher-Order and Symbolic Computation, 21(1– 2):59–88, June 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1388-3690&
                                        volume=21&issue=1&spage=59.

Hu:2008:PED

[177] Zhenjiang Hu, Shin-Cheng Mu, and Masato Takeichi. A programmable editor for developing structured documents based on bidirectional transformations. Higher-Order and Symbolic Computation, 21(1–2):89–118, June 2008. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=21&issue=1&spage=89.

Hulaas:2008:PTL

[178] Jarle Hulaas and Walter Binder. Program transformations for light-weight CPU accounting and control in the Java Virtual Machine. Higher-Order and Symbolic Computation, 21(1– 2):119–146, June 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=21&issue=1&spage=119.

Ochoa:2008:DSL

[179] Claudio Ochoa, Josep Silva, and Germán Vidal. Dynamic slicing of lazy functional programs based on redex trails. Higher-Order and Symbolic Computation, 21(1–2):147–192, June 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1388-3690&volume=21&issue=1&spage=
                                        147.

Pettorossi:2008:TCLa

[180] Alberto Pettorossi and Maurizio Proietti. Totally correct logic program transformations via well-founded annotations. Higher-Order and Symbolic Computation, 21(1–2):193–234, June 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=21&issue=1&spage=
193.

Pettorossi:2008:TCLb

[181] Alberto Pettorossi and Maurizio Proietti. Totally correct logic program transformations via well-founded annotations. Higher-Order and Symbolic Computation, 21(1–2):235, June 2008. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=21&issue=1&spage=235.

Danvy:2008:Ea

[182] Olivier Danvy, R. Kent Dybvig, Julia Lawall, and Peter Thiemann. Editorial. Higher-Order and Symbolic Computation, 21(3):237–238, September 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=21&issue=3&spage=
237.

Skalka:2008:TTE

[183] Christian Skalka. Types and trace effects for object orientation. Higher-Order and Symbolic Computation, 21 (3):239–282, September 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=21&issue=3&spage=239.

GlAck:2008:IJO

[184] Robert Glück. An investigation of Jones optimality and BTI-universal specializers. Higher-Order and Symbolic Computation, 21(3):283–309, September 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=21&issue=3&spage=
283.

Bischof:2008:IAD

[185] Christian H. Bischof, Paul D. Hovland, and Boyana Norris. On the implementation of automatic differentiation tools. Higher-Order and Symbolic Computation, 21(3):311– 331, September 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=21&issue=3&spage=311.

Ziarek:2008:FTS

[186] Lukasz Ziarek, Stephen Weeks, and Suresh Jagannathan. Flattening tuples in an SSA intermediate representation. Higher-Order and Symbolic Computation, 21(3):333–358, September 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=21&issue=3&spage=
333.

Danvy:2008:Eb

[187] Olivier Danvy and Ian Mason. Editorial. Higher-Order and Symbolic Computation, 21(4):359, December 2008. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=21&issue=4&spage=359.

Siskind:2008:NFM

[188] Jeffrey Mark Siskind and Barak A.

Pearlmutter. Nesting forward-mode AD in a functional framework. Higher-Order and Symbolic Computation, 21 (4):361–376, December 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=21&issue=4&spage=361.

Owens:2008:AFP

[189] Scott Owens and Konrad Slind.

Adapting functional programs to higher order logic. Higher-Order and Symbolic Computation, 21(4): 377–409, December 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=21&issue=4&spage=377.

Amtoft:2008:FST

[190] Torben Amtoft. Flow-sensitive type systems and the ambient calculus. Higher-Order and Symbolic Computation, 21 (4):411–442, December 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1388-3690&
                                        volume=21&issue=4&spage=411.

Danvy:2009:Ea

[191] Olivier Danvy and Julian Padget. Editorial. Higher-Order and Symbolic Computation, 22(1):1, March 2009. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=22&issue=1&spage=1.

Hirschowitz:2009:CER

[192] Tom Hirschowitz, Xavier Leroy, and J. B. Wells. Compilation of extended recursion in call-by-value functional languages. Higher-Order and Symbolic Computation, 22(1):3–66, March 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1388-3690&volume=22&issue=1&spage=
                                        3.

Naylor:2009:ESF

[193] Matthew Naylor and Colin Runciman. Expressible sharing for functional circuit description. Higher-Order and Symbolic Computation, 22 (1):67–80, March 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=22&issue=1&spage=67.

Bresson:2009:VLC

[194] Jean Bresson, Carlos Agon, and Gérard Assayag. Visual Lisp/CLOS programming in OpenMusic. Higher-Order and Symbolic Computation, 22 (1):81–111, March 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=22&issue=1&spage=81.

Danvy:2009:Eb

[195] Olivier Danvy and Carolyn L. Talcott.

Editorial. Higher-Order and Symbolic Computation, 22(2):113, June 2009. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=22&issue=2&spage=113.

Stump:2009:DRM

[196] Aaron Stump. Directly reflective meta-programming. Higher-Order and Symbolic Computation, 22 (2):115–144, June 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=22&issue=2&spage=115.

Monniaux:2009:MLW

[197] David Monniaux. A minimalistic look at widening operators. Higher-Order and Symbolic Computation, 22 (2):145–154, June 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=22&issue=2&spage=145.

Johann:2009:PAP

[198] Patricia Johann and Neil Ghani. A principled approach to programming with nested types in haskell. Higher-Order and Symbolic Computation, 22 (2):155–189, June 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=22&issue=2&spage=155.

Danvy:2009:PJL

[199] Olivier Danvy. Peter J. Landin (1930– 2009). Higher-Order and Symbolic Computation, 22(2):191–195, June 2009. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=22&issue=2&spage=191.

Danvy:2009:Ec

[200] Olivier Danvy and Carolyn L. Talcott. Editorial. Higher-Order and Symbolic Computation, 22(3): 197, September 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=22&issue=3&spage=197.

Dargaye:2009:VFH

[201] Zaynah Dargaye and Xavier Leroy. A verified framework for higher-order uncurrying optimizations. Higher-Order and Symbolic Computation, 22 (3):199–231, September 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=22&issue=3&spage=199.

Ariola:2009:TTF

[202] Zena M. Ariola, Hugo Herbelin, and Amr Sabry. A type-theoretic foundation of delimited continuations. Higher-Order and Symbolic Computation, 22 (3):233–273, September 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=22&issue=3&spage=233.

Asai:2009:TDC

[203] Kenichi Asai. On typing delimited continuations: three new solutions to the printf problem. Higher-Order and Symbolic Computation, 22 (3):275–291, September 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=22&issue=3&spage=275.

Danvy:2009:E

[204] Olivier Danvy and Carolyn L. Talcott. Editorial. Higher-Order and Symbolic Computation, 22(4): 293, December 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=22&issue=4&spage=293.

Bornat:2009:PLA

[205] Richard Bornat. Peter Landin:

a computer scientist who inspired a generation. Higher-Order and Symbolic Computation, 22(4):295– 298, December 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=22&issue=4&spage=295.

Boca:2009:PRP

[206] Paul P. Boca. Personal recollections of Peter Landin: 1987–2009. Higher-Order and Symbolic Computation, 22 (4):299–303, December 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=22&issue=4&spage=299.

Hammond:2009:PLP

[207] Kevin Hammond and Greg Michaelson. The Peter Landin Prize. Higher-Order and Symbolic Computation, 22 (4):305–312, December 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=22&issue=4&spage=305.

Tennent:2009:ILS

[208] R. D. Tennent. An introduction to Landin’s “Getting Rid of Labels”. Higher-Order and Symbolic Computation, 22(4):313–314, December 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1388-3690&volume=22&issue=4&spage=
                                        313.

Landin:2009:GRL

[209] P. J. Landin. Getting rid of labels. Higher-Order and Symbolic Computation, 22(4):315–329, December 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=22&issue=4&spage=

315. Introduced by Bob Tennent.

Clark:2009:SAC

[210] Tony Clark. Stories about calculations:

remembering Peter Landin. Higher-Order and Symbolic Computation, 22 (4):331–332, December 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=22&issue=4&spage=331.

Landin:2009:CHH

[211] Peter J. Landin. Calculations: a hole in the heart of the study of computing. Higher-Order and Symbolic Computation, 22(4):333– 359, December 2009. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=22&issue=4&spage=333. Intro-

duced by Tony Clark.

Alves:2010:LIT

[212] Sandra Alves, Maribel Fernández, Mário Florido, and Ian Mackie. Linearity and iterator types for Gödel’s system. Higher-Order and Symbolic Computation, 23(1):1–27, March 2010. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=23&issue=1&spage=1.

Spoto:2010:MSL

[213] Fausto Spoto and Étienne Payet. Magicsets for localised analysis of Java bytecode. Higher-Order and Symbolic Computation, 23(1):29–86, March 2010. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=23&issue=1&spage=29.

Simon:2010:TVP

[214] Axel Simon and Andy King. The two variable per inequality abstract domain. Higher-Order and Symbolic Computation, 23(1):87–143, March 2010. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=23&issue=1&spage=87.

Middelkoop:2010:LSG

[215] Arie Middelkoop, Atze Dijkstra, and S. Doaitse Swierstra. A lean specification for GADTs: system F with first-class equality proofs. Higher-Order and Symbolic Computation, 23 (2):145–166, June 2010. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=23&issue=2&spage=145.

Herman:2010:SEG

[216] David Herman, Aaron Tomb, and Cormac Flanagan. Space-efficient gradual typing. Higher-Order and Symbolic Computation, 23(2):167–189, June 2010. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=23&issue=2&spage=
167.

Grov:2010:HBC

[217] Gudmund Grov and Greg Michaelson.

Hume box calculus: robust system development through software transformation. Higher-Order and Symbolic Computation, 23(2):191–226, June 2010. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=23&issue=2&spage=191.

Sculthorpe:2010:KCF

[218] Neil Sculthorpe and Henrik Nilsson. Keeping calm in the face of change: Towards optimisation of FRP by reasoning about change. Higher-Order and Symbolic Computation, 23 (2):227–271, June 2010. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=23&issue=2&spage=227.

Gallagher:2010:E

[219] John P. Gallagher and Janis Voigtländer.

Editorial. Higher-Order and Symbolic Computation, 23(3):273–274, September 2010. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=23&issue=3&spage=
273.

Lakhotia:2010:CSA

[220] Arun Lakhotia, Davidson R. Boccardo, Anshuman Singh, and Aleardo Manacero. Context-sensitive analysis without calling-context. Higher-Order and Symbolic Computation, 23 (3):275–313, September 2010. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=23&issue=3&spage=275.

Holdermans:2010:MSM

[221] Stefan Holdermans and Jurriaan Hage. Making “stricterness” more relevant. Higher-Order and Symbolic Computation, 23(3):315–335, September 2010. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=23&issue=3&spage=
315.

Henglein:2010:GMP

[222] Fritz Henglein and Ken Friis Larsen. Generic multiset programming with discrimination-based joins and symbolic Cartesian products. Higher-Order and Symbolic Computation, 23 (3):337–370, September 2010. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=23&issue=3&spage=337.

Rudolph:2010:MTS

[223] Johannes Rudolph and Peter Thiemann. Mnemonics: type-safe bytecode generation at run time. Higher-Order and Symbolic Computation, 23 (3):371–407, September 2010. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=23&issue=3&spage=371.

Serrano:2010:MTS

[224] Manuel Serrano and Christian Queinnec. A multi-tier semantics for Hop. Higher-Order and Symbolic Computation, 23(4):409–431, November 2010. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=23&issue=4&spage=
409.

Chadwick:2010:FAG

[225] Bryan Chadwick and Karl Lieberherr. A functional approach to generic programming using adaptive traversals. Higher-Order and Symbolic Computation, 23 (4):433–463, November 2010. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=23&issue=4&spage=433.

Achten:2010:TTF

[226] Peter Achten, Marko van Eekelen, Pieter Koopman, and Marco T. Morazán. Trends in Trends in Functional Programming 1999/2000 versus 2007/2008. Higher-Order and Symbolic Computation, 23(4): 465–487, November 2010. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=23&issue=4&spage=465.

Swierstra:2010:MDT

[227] Wouter Swierstra. More dependent types for distributed arrays. Higher-Order and Symbolic Computation, 23 (4):489–506, November 2010. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=23&issue=4&spage=489.

Lawall:2011:E

[228] Julia Lawall, Germán Puebla, and Germán Vidal. Editorial. Higher-Order and Symbolic Computation, 24 (1–2):1–2, June 2011. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=24&issue=1&spage=1.

elBoustani:2011:ITE

[229] Nabil el Boustani and Jurriaan Hage. Improving type error messages for generic Java. Higher-Order and Symbolic Computation, 24(1–2):3–39, June 2011. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1388-3690&volume=24&issue=1&spage=
                                        3.

Simmons:2011:LAP

[230] Robert J. Simmons and Frank Pfenning. Logical approximation for program analysis. Higher-Order and Symbolic Computation, 24(1–2):41–80, June 2011. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=1388-3690&
volume=24&issue=1&spage=41.

Salama:2011:SCC

[231] Cherif Salama, Gregory Malecha, Walid Taha, Jim Grundy, and John O’Leary. Static consistency checking for Verilog wire interconnects: Using dependent types to check the sanity of Verilog descriptions. Higher-Order and Symbolic Computation, 24 (1–2):81–114, June 2011. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=24&issue=1&spage=81.

Pardo:2011:SFR

[232] Alberto Pardo, João Paulo Fernandes, and João Saraiva. Shortcut fusion rules for the derivation of circular and higher-order programs. Higher-Order and Symbolic Computation, 24 (1–2):115–149, June 2011. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=24&issue=1&spage=115.

Chen:2011:SEL

[233] Kung Chen, Shu-Chun Weng, Jia-Yin Lin, Meng Wang, and Siau-Cheng Khoo. Side-effect localization for lazy, purely functional languages via aspects. Higher-Order and Symbolic Computation, 24(1–2):151–189, June 2011. CO-DEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=24&issue=1&spage=151.

Niqui:2011:PMT

[234] M. Niqui and J. J. M. M. Rutten. A proof of Moessner’s theorem by coinduction. Higher-Order and Symbolic Computation, 24(3): 191–206, September 2011. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=24&issue=3&spage=191.

Garrigue:2011:PRN

[235] Jacques Garrigue and Keiko Nakata. Path resolution for nested recursive modules. Higher-Order and Symbolic Computation, 24(3):207– 237, September 2011. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1388-3690&
                                        volume=24&issue=3&spage=207.

Kerneis:2011:CPC

[236] Gabriel Kerneis and Juliusz Chroboczek. Continuation-Passing C, compiling threads to events through continuations. Higher-Order and Symbolic Computation, 24(3):239–279, September 2011. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1388-3690&volume=24&issue=3&spage=

239. See erratum [240].

Iwasaki:2011:PIS

[237] Hideya Iwasaki, Takeshi Morimoto, and Yasunao Takano. Pruning with improving sequences in lazy functional programs. Higher-Order and Symbolic Computation, 24(4): 281–309, November 2011. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=24&issue=4&spage=281.

Asai:2011:FU

[238] Kenichi Asai, Oleg Kiselyov, and Chung chieh Shan. Functional un|unparsing. Higher-Order and Symbolic Computation, 24(4):311– 340, November 2011. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1388-3690&
volume=24&issue=4&spage=311.

Liu:2011:TSS

[239] Yu David Liu, Christian Skalka, and Scott F. Smith. Type-specialized staged programming with process separation. Higher-Order and Symbolic Computation, 24(4):341–385, November 2011. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1388-3690&volume=24&issue=4&spage=
341.

Kerneis:2011:ECP

[240] Gabriel Kerneis and Juliusz Chroboczek.

Erratum to: Continuation–Passing C, compiling threads to events through continuations. Higher-Order and Symbolic Computation, 24(4):387, November 2011. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=

                                        1388-3690&volume=24&issue=4&spage=

387. See [236].

Kiselyov:2012:EP

[241] Oleg Kiselyov, Julia Lawall, and Simon Thompson. Editorial PEPM2012. Higher-Order and Symbolic Computation, 25(1):1–2, March 2012. CODEN LSCOEX. ISSN 1388-3690 (print), 1573-0557 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10990-013-9092-0;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s10990-013-9092-0.pdf.

Matsuda:2012:PTI

[242] Kazutaka Matsuda, Kazuhiro Inaba, and Keisuke Nakano. Polynomial-time inverse computation for accumulative functions with multiple data traversals. Higher-Order and Symbolic Computation, 25(1):3–38, March 2012. CO-DEN LSCOEX. ISSN 1388-3690 (print), 1573-0557 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10990-013-9097-8;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s10990-013-9097-8.pdf.

Kobayashi:2012:FPC

[243] Naoki Kobayashi, Kazutaka Matsuda, and Ayumi Shinohara. Functional programs as compressed data. Higher-Order and Symbolic Computation, 25(1):39–84, March 2012. CO-DEN LSCOEX. ISSN 1388-3690 (print), 1573-0557 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10990-013-9093-z.

Degen:2012:ICL

[244] Markus Degen, Peter Thiemann, and Stefan Wehr. The interaction of contracts and laziness. Higher-Order and Symbolic Computation, 25(1):85– 125, March 2012. CODEN LSCOEX. ISSN 1388-3690 (print), 1573-0557 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10990-013-
9094-y.

Sasano:2012:ACV

[245] Isao Sasano and Takumi Goto. An approach to completing variable names for implicitly typed functional languages. Higher-Order and Symbolic Computation, 25(1):127–163, March 2012. CODEN LSCOEX. ISSN 1388-3690 (print), 1573-0557 (electronic). URL

http://link.springer.com/article/
10.1007/s10990-013-9095-x.

Rompf:2012:SVL

[246] Tiark Rompf, Nada Amin, and Adriaan Moors. Scala-Virtualized: linguistic reuse for deep embeddings. Higher-Order and Symbolic Computation, 25(1):165–207, March 2012. CODEN LSCOEX. ISSN 1388-3690 (print), 1573-0557 (electronic). URL

http://link.springer.com/article/
10.1007/s10990-013-9096-9.

Klein:2012:RVM

[247] Casey Klein, Matthew Flatt, and Robert Bruce Findler. The Racket virtual machine and randomized testing. Higher-Order and Symbolic Computation, 25(2–4):209–253, December 2012. CODEN LSCOEX. ISSN 1388-3690 (print), 1573-0557 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10990-013-9091-1;   http://link.
springer.com/article/10.1007/s10990-
013-9091-1.

Gill:2012:TAT

[248] Andy Gill, Tristan Bull, and Andrew Farmer. Types and associated type families for hardware simulation and synthesis. Higher-Order and Symbolic Computation, 25(2–4):255–274, December 2012. CODEN LSCOEX. ISSN 1388-3690 (print), 1573-0557 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10990-013-9098-7;   http://link.
                                        springer.com/article/10.1007/s10990-
                                        013-9098-7.

Capper:2012:STS

[249] John Capper and Henrik Nilsson. Structural types for systems of equations. Higher-Order and Symbolic Computation, 25(2–4):275–310, December 2012. CODEN LSCOEX. ISSN 1388-3690 (print), 1573-0557 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10990-013-9099-6;   http://link.
                                        springer.com/article/10.1007/s10990-
                                        013-9099-6.

Zsok:2013:ET

[250] Viktória Zsók, Rex Page, and Julia Lawall. Editorial TFP 2009/2010. Higher-Order and Symbolic Computation, 26(1–4):1–2, December 2013. CO-DEN LSCOEX. ISSN 1388-3690 (print), 1573-0557 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10990-014-9102-x;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s10990-014-9102-x.pdf.

Horstmeyer:2013:GBC

[251] Thomas Horstmeyer and Rita Loogen. Graph-based communication in Eden.

Higher-Order and Symbolic Computation, 26(1–4):3–28, December 2013. CODEN LSCOEX. ISSN 1388-3690 (print), 1573-0557 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10990-014-9101-y;   http://link.
springer.com/article/10.1007/s10990-
014-9101-y.

Clerici:2013:GIT

[252] Silvia Clerici and Cristina Zoltan.

Graphical and incremental type inference. A graph transformation approach. Higher-Order and Symbolic Computation, 26(1–4):29–62, December 2013. CODEN LSCOEX. ISSN 1388-3690 (print), 1573-0557 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10990-014-9104-8;   http://link.
springer.com/article/10.1007/s10990-
014-9104-8.

Zerny:2013:GRR

[253] Ian Zerny. On graph rewriting, reduction, and evaluation in the presence of cycles. Higher-Order and Symbolic Computation, 26(1–4):63–84, December 2013. CODEN LSCOEX. ISSN 1388-3690 (print), 1573-0557 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10990-014-9103-9;   http://link.
springer.com/article/10.1007/s10990-
014-9103-9.